Saving, Investment, and the Financial System (Mobile-Enabled) Suppose GDP in this country is $1,175 milion. Enter the amount for investment. Value Consumption (C)
media%2Fb3e%2Fb3e6fc72-04f6-4b09-ac78-f1
media%2F989%2F98919401-9d9a-45d9-9ed2-01
media%2F8e1%2F8e1126a8-c1c0-4c1e-a731-6d
media%2Fd86%2Fd8699842-8c63-462b-8303-31
media%2F555%2F55532b6b-ce70-49b6-9dd8-fb
Answer #1

(a) National saving ($Million) = (Y - C - T) + (T - G) = Y - C - G = I = 1,175 - 625 - 250 = 300

(b) Private saving ($Million) = Y - C - T = 1,175 - 625 - 225 = 325

(c) Public saving ($Million) = T - G = 225 - 250 = - 25

(d) Since T < G, Government is running a budget deficit.